perpendicular to 5x + 4y = -52 through (8,4)
Answer by nerdybill(7384)

perpendicular to 5x + 4y = -52 through (8,4)
.
First, determine the slope of:
5x + 4y = -52
Do this by changing to "slope-intercept" form:
4y = -5x-52
y = (-5/4)x - 13
Slope is -5/4
.
For a line to be perpendicular, it must be a "negative reciprocal":
m(-5/4) = -1
m = (-1)(-4/5)
m = 4/5
.
So, our NEW line must have
slope of 4/5
crossing (8,4)
.
Plug the above into the "point-slope" form:
y-y1 = m(x-x1)
y-4 = (4/5)(x-8)
y-4 = (4/5)x - (4/5)8
y-4 = (4/5)x - 32/5
y = (4/5)x - 32/5 + 4
y = (4/5)x - 32/5 + 20/5
y = (4/5)x - 12/5 (this is what they're looking for)